# Leveraging Online Data to Enhance Medical Knowledge in a Small Persian Language Model

### v3

@@ Abstract
The rapid advancement of language models has demonstrated the potential of artificial intelligence in the healthcare industry. However, small language models struggle with specialized domains in low-resource languages like Persian. While numerous medical-domain websites exist in Persian, no curated dataset or corpus has been available—making ours the first of its kind. This study explores the enhancement of medical knowledge in a small language model by leveraging accessible online data, including a crawled corpus from medical magazines and a dataset of real doctor-patient Q&A pairs.
We fine-tuned a baseline model using our curated data to improve its medical knowledge. Benchmark evaluations demonstrate that the fine-tuned model achieves improved accuracy in medical question answering and provides better responses compared to its baseline. Notably, the trained model successfully passed the Iranian Basic Medical Science Entrance Exam, taken in September 2023, and improved Persian-translated MMLU accuracy by an average of 2.67%. This work highlights the potential of leveraging open-access online data to enrich small language models in medical fields, providing a novel solution for Persian medical AI applications suitable for resource-constrained environments.
@@ I Introduction
Moreover, the development of small language models is particularly crucial in the medical domain due to privacy concerns. These models can be optimized to run on local devices, ensuring that sensitive patient data remains secure and confidential, which is a paramount consideration in healthcare settings. However, the unavailability of appropriate medical corpora and datasets in Persian has hindered progress in this area, impeding the creation of robust language models that can effectively address the linguistic and cultural nuances of the Persian-speaking population.
@@
  • Introducing the first open-source 222we have published our model and datasets at https://huggingface.co/gaokerena persian medical language model that achieved state of the art result in comparison to other home device runnable alternatives
@@ III Baseline Model
We have chosen aya-expanse-8b as our baseline model primarily due to the lack of open-source Persian medical language models, which necessitates the use of a general-purpose language model. While there are several multilingual options available, including aya-expanse [5], gemma-2 [15], qwen2.5 [16], and persianmind [17], we have determined that aya-expanse is the most suitable choice for our needs. One key reason is that the training data for the other models predominantly consists of non-Persian languages, leading to biases that may result in the generation of non-Persian characters, even when we explicitly instruct the model to use only Persian. In contrast, aya-expanse demonstrates a robust understanding of Persian grammar and produces grammatically rich Persian text, making it a more reliable option for our research. Additionally, we conducted an ablation study by fine-tuning qwen2.5-7b, and the results further validated our decision to select aya-expanse-8b as the baseline model. For more details, please refer to the ablation study section. Furthermore, if we merge our updated parameters into aya-vision [18], another model from the aya family, instead of aya-expanse, we gain the capability to incorporate medical images such as MRIs and CT scans as inputs, thereby enhancing our model’s ability to process complex medical questions.
@@ IV-A Corpus
Figure 1: Our corpus resources
@@ IV-B Dataset
Figure 2: MF3QA resources
@@ V-A Fine tuning
We fine-tuned the 8-billion-parameter checkpoint of the aya-expanse model on 60% of our corpus, focusing on minimizing resource usage. To ensure an efficient fine-tuning process, we employed gradient checkpointing and a small batch size of 2,reducing memory requirements during training. Additionally,we used gradient accumulation steps of 16, effectively increasing the overall batch size to 32 and enabling stable training dynamics. To further reduce the memory usage of our fine-tuning process, we leveraged Low Rank Adaptation (LoRA) [28] to significantly reduce the number of trainable parameters. Specifically, we implemented a rank of 8, an alpha value of 16, and a dropout rate of 5%, applying the LoRA weights to all trainable parameters of each transformer layer.To further optimize the process, we employed efficient tokenization and memory-aware training techniques. The tokenization process divided the input text into manageable token sequences, ensuring consistent input and label structures by truncating, padding, and handling overflowed tokens to maintain contextual integrity within a fixed context length. This streamlined preparation, coupled with the LoRA-based fine-tuning, was further enhanced by Flash Attention 2 [29].By minimizing memory overhead, Flash Attention 2 allowed us to handle longer context lengths and larger batch sizes efficiently, enabling effective fine-tuning for the next-token prediction objective while balancing computational efficiency with model performance. Figure 3 presents the loss curve during the fine-tuning stage. The curve demonstrates a smooth and consistent decrease in training loss, indicating stable convergence without signs of divergence or severe overfitting.
@@ V-C Carbon Footprint
The carbon footprint of our model optimization—including both fine-tuning and instruction-tuning—was estimated based on hardware specifications and operational duration. The process ran for a combined total of 19 hours on NVIDIA A100 PCIe 40GB GPUs hosted in Google Cloud Platform’s asia-east1 region, utilizing a Google Colab Pro+ account for the hardware resources. Assuming a typical power consumption of 250 watts per GPU, the total energy used was 4.75 KWh (250 watts × 19 hours). Using the carbon intensity factor of the asia-east1 grid (0.56 kilograms of CO2 equivalent per kWh), this translates to 2.66 kilograms of CO2 equivalent emitted during the tuning process [30].
@@ VI-A Comparison with general purpose language models
As you can see in Table V, our model achieved remarkable success by surpassing the passing score of the Iranian Basic Medical Sciences Entrance Exam, which stands at 36%, making it the first Persian language model with fewer than 8 billion parameters to pass this exam. Furthermore, our model demonstrated improvements on the translated MMLU dataset, achieving not only higher average scores but also excelling across most sub-categories, thereby showcasing its effectiveness in understanding and generating medical knowledge in the Persian language. Notably, the reason PersianMind has a low inference time, as shown in Table V, is that it tends to generate very short responses compared to other models, producing the end-of-sentence token much sooner. In addition to our multiple-choice question-answering evaluation, we also utilized GPT-4o [32] as an evaluator for free-form question answering. We provided the test set from the MF3QA dataset (which includes 2,000 records) to both the competing language model and our model. As shown in Figure 5, GPT-4o predominantly preferred the responses generated by our model over those from the other three language models tested. This indicates that our model delivers high-quality responses, as judged by an advanced language model. Moreover, GPT-4o was prompted to prefer responses that provide more accurate medical information while avoiding incorrect medical information, further validating the reliability and precision of our model in generating medically accurate answers. 555you can see the prompt at https://github.com/Mehrdadghassabi/Gaokerena-V/blob/main/evaluation/free_form_qa/gaokerena_vs_aya/Untitled20.ipynb
@@ VI-B Comparison with pipeline alternatives
VII Ablation study
As an ablation study, we trained qwen2.5 using exactly the same methodology and training setup as employed for aya-expanse. This allowed us to directly compare the performance of the two baseline models under identical conditions, the results of this comparison are presented in Figure 7 and detailed further in Table VII, where it is evident that the trained qwen-2.5 model achieved lower scores across key performance metrics in comparison with gaokerena-V, which utilized Aya-Expanse as its baseline model. These findings highlight the advantages of using aya-expanse as the backbone model and reinforce its effectiveness in producing higher-performing models.
TABLE VII: our model performance in comparison with pipeline alternatives
gaokerena-V Qwen2.5 Qwen2.5
(ours) trained
MMLU-
anatomy(fa) 48.14 41.48 36.29
MMLU-
medical-genetics(fa) 53.0 52.0 50.0
MMLU-
college-medicine(fa) 43.93 43.35 42.77
MMLU-
clinical-knowledge(fa) 55.47 47.92 49.81
MMLU-
professional-medicine(fa) 47.05 43.01 42.27
MMLU-
college-biology(fa) 47.22 42.36 0.50
MMLU(avg) 49.31 45.17 45.35
IBMSEE Sept2023 38.69 33.33 35.71
Number of
parameters 8b 7.6b 7.6b
inference time $\approx 10s$ $\approx 15s$ $\approx 15s$
Figure 7: Our model win rate against trained qwen2.5
@@ VII Ablation study
As mentioned in the baseline model section, we can merge our updated parameters into aya-vision instead of aya-expanse, This will allow us to incorporate medical images, such as MRIs and CT scans, as inputs to our model. However, it is important to note that our models capabilities with multi-modal queries are under-explored. Our future research plans include exploring and improving the model’s responses to queries accompanied by medical images. This will involve fine-tuning the model to better process and align textual and visual data, with the goal of expanding its utility in medical diagnostics and analysis.

### v5

@@ Abstract
The rapid advancement of language models has demonstrated the potential of artificial intelligence in the healthcare industry. However, small language models struggle with specialized domains in low-resource languages like Persian. While numerous medical-domain websites exist in Persian, no curated dataset or corpus has been available—making ours the first of its kind. This study introduces a newly curated dataset comprising 20k doctor-patient Q&A pairs and 60% of a 90-million-token crawled corpus from medical magazines. Using a parameter-efficient fine-tuning approach, we enhanced the medical knowledge of the baseline model, aya-expanse-8b. Benchmark evaluations demonstrate that the fine-tuned model achieves improved accuracy in medical question answering and successfully passed the Iranian Basic Medical Science Entrance Exam (IBSEE) in September 2023, which the baseline model did not. Additionally, the fine-tuned model improved Persian-translated MMLU accuracy by an average of 2.67%. This work highlights the potential of leveraging open-access online data to enrich small language models in medical fields, providing a novel solution for Persian medical AI applications suitable for resource-constrained environments. Future research could explore multimodal input to further enhance performance.
@@ I Introduction
Moreover, the development of small language models is particularly crucial in the medical domain due to privacy concerns. These models can be optimized to run on local devices, ensuring that sensitive patient data remains secure and confidential, which is a paramount consideration in healthcare settings [5]. However, the unavailability of appropriate medical corpora and datasets in Persian has hindered progress in this area, impeding the creation of robust language models that can effectively address the linguistic and cultural nuances of the Persian-speaking population.
@@
  • Introducing the first open-source 222we have published our model and datasets at https://huggingface.co/gaokerena persian medical language model that achieved promissing result in comparison to other home device runnable alternatives
@@ III Baseline Model
We have chosen aya-expanse-8b as our baseline model primarily due to the lack of open-source Persian medical language models, which necessitates the use of a general-purpose language model. While there are several multilingual options available, including aya-expanse, gemma-2 [16], qwen2.5 [17], and persianmind [18], we have determined that aya-expanse is the most suitable choice for our needs. One key reason is that the training data for the other models predominantly consists of non-Persian languages, leading to biases that may result in the generation of non-Persian characters, even when we explicitly instruct the model to use only Persian. In contrast, aya-expanse demonstrates a robust understanding of Persian grammar and produces grammatically rich Persian text, making it a more reliable option for our research. Furthermore, if we merge our updated parameters into aya-vision [19], another model from the aya family, instead of aya-expanse, we gain the capability to incorporate medical images such as MRIs and CT scans as inputs, thereby enhancing our model’s ability to process complex medical questions.
@@ IV-A Corpus
Figure 1: Percentage of content from each magazine crawled for the corpus
@@ IV-B Dataset
Figure 2: Percentage of content from each forums crawled for MF3QA dataset
@@ V-A Fine tuning
We fine-tuned the 8-billion-parameter checkpoint of the aya-expanse model on 60% of our corpus, focusing on minimizing resource usage.This decision was driven by the need to decrease the cost of training, as this project was entirely self-funded and constrained by a limited budget. To ensure an efficient fine-tuning process, we employed gradient checkpointing and a small batch size of 2,reducing memory requirements during training. Additionally,we used gradient accumulation steps of 16, effectively increasing the overall batch size to 32 and enabling stable training dynamics.
To further reduce the memory usage of our fine-tuning process, we leveraged Low Rank Adaptation (LoRA) [29] to significantly reduce the number of trainable parameters. Specifically, we implemented a rank of 8, an alpha value of 16, and a dropout rate of 5%, applying the LoRA weights to all trainable parameters of each transformer layer.To further optimize the process, we employed efficient tokenization and memory-aware training techniques. The tokenization process divided the input text into manageable token sequences, ensuring consistent input and label structures by truncating, padding, and handling overflowed tokens to maintain contextual integrity within a fixed context length.
This streamlined preparation, coupled with the LoRA-based fine-tuning, was further enhanced by Flash Attention 2 [30].By minimizing memory overhead, Flash Attention 2 allowed us to handle longer context lengths and larger batch sizes efficiently, enabling effective fine-tuning for the next-token prediction objective while balancing computational efficiency with model performance. Figure 3 presents the loss curve during the fine-tuning stage. The curve demonstrates a smooth and consistent decrease in training loss, indicating stable convergence without signs of divergence or severe overfitting.
@@ V-C Carbon Footprint
The carbon footprint of our model optimization—including both fine-tuning and instruction-tuning—was estimated based on hardware specifications and operational duration. The process ran for a combined total of 19 hours on a NVIDIA A100 PCIe 40GB GPU hosted in Google Cloud Platform’s asia-east1 region, utilizing a Google Colab Pro+ account for the hardware resources. Assuming a typical power consumption of 250 watts per GPU, the total energy used was 4.75 KWh (250 watts × 19 hours). Using the carbon intensity factor of the asia-east1 grid (0.56 kilograms of CO2 equivalent per kWh), this translates to 2.66 kilograms of CO2 equivalent emitted during the tuning process [31].
@@ VI Results
Notice that all experiments were conducted on an NVIDIA L4 GPU, which is available through the Google Colab Pro and Pro+ plans. Leveraging this hardware ensured that the experiments were carried out in an accessible yet capable computational environment, reflecting practical limitations similar to those faced by many researchers.
@@ VI-A Comparison with general purpose language models
As shown in Table V, our model achieved remarkable success by surpassing the passing score of the Iranian Basic Medical Sciences Entrance Exam, which is set at 36%. This achievement makes it the first Persian language model with fewer than 8 billion parameters to pass this exam. Furthermore, our model demonstrated significant improvements on the translated MMLU dataset, achieving higher average scores and excelling across most sub-categories. This highlights its effectiveness in understanding and generating medical knowledge in the Persian language.
Notably, in the anatomy category, the model achieved a 7.4% improvement compared to the baseline model. This improvement is likely due to the availability of more anatomy-related data in authoritative online magazines compared to other categories. However, it also underscores the lack of publicly available medical data in the Persian language, which poses challenges for further enhancing the model’s performance across other medical subcategories.
Another key observation, as shown in Table V, is that PersianMind exhibits low inference time, primarily because it generates very short responses compared to other models, producing the end-of-sentence token much sooner.
In addition to our multiple-choice question-answering evaluation, we utilized GPT-4o [33] as an evaluator for free-form question answering. We provided the test set from the MF3QA dataset, which includes 2,000 records, to both our model and competing language models. As shown in Figure 5, GPT-4o predominantly preferred the responses generated by our model over those from the other three language models tested. This demonstrates that our model delivers high-quality responses, as judged by an advanced language model.
Moreover, GPT-4o was prompted to prefer responses that provide more accurate medical information while avoiding incorrect medical information. This further validates the reliability and precision of our model in generating medically accurate answers. 555you can see the prompt at https://github.com/Mehrdadghassabi/Gaokerena-V/blob/main/evaluation/free_form_qa/gaokerena_vs_aya/Untitled20.ipynb
@@ VI-B Comparison with pipeline alternatives
Although Gaokerena-V demonstrates enhanced Persian medical knowledge compared to its baseline and other alternatives, its current performance is not yet sufficient for reliable use in clinical scenarios. As mentioned in the baseline model section, we plan to merge our updated parameters into aya-vision rather than aya-expanse, enabling the model to incorporate multimodal inputs such as medical images (e.g., MRIs and CT scans). However, the model’s capabilities in handling multimodal queries remain under-explored. Our research will focus on improving its ability to interpret and align textual and visual information, with the goal of expanding its diagnostic and analytical utility in medicine. Additionally, future work will involve collaboration with real doctors to enhance the model through reinforcement learning from human feedback (RLHF) [36] and rigorous clinical validation, ensuring its practical applicability and safety in real-world healthcare settings.